OpenStudy (anonymous):

Given the function f(x) = 2(x − 1)^2 + 1, indicate the shifts that will affect the location of the vertex and explain what effect they will have. Use complete sentences. f(x+3) f(x) + 3 f(3x) 3•f(x)

OpenStudy (anonymous):

@phi

OpenStudy (phi):

what is the vertex of f(x) = 2(x − 1)^2 + 1, ?

OpenStudy (anonymous):

the axis of symmetry is x = 4 I believe because the x of the vertex is 4

OpenStudy (anonymous):

oh wait wrong question lol, working on a couple at once

OpenStudy (phi):

your equation is in "vertex form" y = a(x-h)^2 + k (h,k) is the vertex

OpenStudy (anonymous):

alright the vertex is 1,1 I think

OpenStudy (phi):

yes, if you replace x with x+3 in 2(x − 1)^2 + 1, and simplify what do you get ?

OpenStudy (anonymous):

2(3 - 1)^2 + 1 distribute... 6 - 2 3 3^3 9 + 1 10?

OpenStudy (anonymous):

3^2, not

OpenStudy (anonymous):

3^3

OpenStudy (phi):

no, we keep the x replace x with x+3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

2(x − 1)^2 + 1 2(x - 3)^2 + 1 distribute... 2x - 6 = -4x -4x^2 = 16 + 1 17

OpenStudy (phi):

we will not get a number, just a different equation \[ 2(x − 1)^2 + 1 \] replace x with (x+3) \[ 2(x+3 − 1)^2 + 1 \] simplify \[ 2(x +2 )^2 + 1 \\ 2(x - (-2) )^2 + 1 \] what is the vertex of this parabola ?

OpenStudy (anonymous):

-2, 1

OpenStudy (phi):

so we moved from (1,1) to (-2,1) it has the same shape because the 2 out front (the "a" parameter) did not change. how would you describe the change ?

OpenStudy (anonymous):

the graph changed by moving to the left 3 units

OpenStudy (phi):

Yes. in general f(x PLUS N) shifts to the LEFT N steps. (counter intuitive or backwards)

OpenStudy (phi):

now do f(x) + 3 where f(x) = 2(x − 1)^2 + 1

OpenStudy (anonymous):

we would make (x) = 2(x − 1)2 + 1 into (x) = 2(x − 1)2 + 3 right?

OpenStudy (phi):

almost. you do not change 1 into 3. you add 3 to what you started with.

OpenStudy (anonymous):

right... (x) = 2(x − 1)2 + 1 + 3? so 2(x - 1)^2 + 4?

OpenStudy (anonymous):

which would take the cordinates from 1,1 to 1,4?

OpenStudy (phi):

yes, and how does the graph change ?

OpenStudy (anonymous):

It moves upwards 3 units

OpenStudy (phi):

yes, the rule is f(x)+N shifts up N steps. (this one is easier to remember)

OpenStudy (phi):

For f(3x) replace x with 3x \[ f(x) = 2(x − 1)^2 + 1 \\ f(x) = 2(3x-1)^2 + 1 \\ f(x) = 2\cdot \sqrt{3}(x-\frac{1}{3})^2 + 1 \] now two things have changed. the vertex shifted (how ?) and the "a" got bigger. A bigger "a" (number out front) makes the parabola "zoom up" faster

OpenStudy (anonymous):

uhh I'm a bit confused on this one..

OpenStudy (anonymous):

the graph is now at 1/3, 1 cordinates??

OpenStudy (phi):

Here it is

OpenStudy (phi):

yes, the vertex is at ⅓,1

OpenStudy (anonymous):

I see so now the spread is much thinner

OpenStudy (phi):

it goes up faster, which makes it narrower. if we changed a to ¼, the zoom up would be slower, and the parabola would be "fatter" or "flatter"

OpenStudy (anonymous):

ok i gotcha

OpenStudy (anonymous):

last one, 3•f(x)

OpenStudy (phi):

start with f(x) = 2(x − 1)^2 + 1 multiply both sides by 3

OpenStudy (anonymous):

so... 3f(x) = 5(x-1)^2 + 1?

OpenStudy (anonymous):

err not 5... 6

OpenStudy (anonymous):

3f(x) = 6(x-1)^2 + 1?

OpenStudy (phi):

you multiply *everything* by 3, not just your favorite terms.

OpenStudy (anonymous):

ok.. 3f(x) = 6(x - 3)^2 + 3

OpenStudy (phi):

example, with numbers: 5 = 4+1 3*5 = 3*1 + 3*4 check: 3*5 =15 and 3*1 + 3*4= 3+12= 15 if we just did 3*5 = 3*4 + 1, it does not work.

OpenStudy (phi):

3f(x) = 6(x - 3)^2 + 3 what happened to the vertex ? and how does the shape change ?

OpenStudy (anonymous):

the vertex moves from 1,1 to 3,3

OpenStudy (anonymous):

so up two units and to the right 2 units

OpenStudy (phi):

oops, wait. f(x) = 2(x − 1)^2 + 1 3f(x) = 3*(2(x − 1)^2 + 1 ) try again

OpenStudy (phi):

you should get 3*2*(x-1)^2 + 3

OpenStudy (anonymous):

so then 6(x-1)^2 + 3?

OpenStudy (phi):

much better

OpenStudy (anonymous):

so the vertex does not move?

OpenStudy (phi):

the original f(x) = 2(x − 1)^2 + 1 has a vertex at (1,1)

OpenStudy (anonymous):

oh wait 1,3

OpenStudy (anonymous):

so the vertex goes up 2 units

OpenStudy (phi):

and the shape ?

OpenStudy (anonymous):

uhh skinnier?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thanks phi! I have a few more qestions that are shorter... Could you help me with those?

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating.